At many established Forward Operating Bases (FOBs)—including the one in this scenario—civilian contracted companies or other organizations provide commercial SIPR services to every unit on the FOB. This can only occur if both the cable infrastructure is proper and the brigade has the required device images on all network devices (this ensures all necessary patches and security measures are present). The probability of proper cable infrastructure is 0.80 and the probability of the required device images is 0.93. The probability of the brigade having the required device images and not having proper cable infrastructure is 0.18.

1) Calculate the probability the Brigade TOC receives commercial SIPR services.

2) Are proper cable infrastructure and brigade device imaging independent events? Justify your answer mathematically.

3) Do you think proper cable infrastructure and brigade device imaging should be independent of each other? Explain your answer.
Answer #1
Cable Infrastructure
Yes No Total
Device Images Yes 0.93 - 0.18 = 0.75 0.18 0.93
No 0.07 - 0.02 = 0.05 0.2 - 0.18 = 0.02 1 - 0.93 = 0.07
Total 0.8 1 - 0.8 = 0.2 1

(1) Probability  = 0.75

(2) P(Proper cable) = 0.8, P(Device imaging) = 0.93

P(Proper cable | Device imaging) = P(Proper cable and Device imaging) / P(Device imaging) = 0.75/0.93 = 0.806

P(Proper cable | Device imaging) ≠ P(Proper cable), which means the two events are not independent

(3) Yes, they should be independent of each other. The basic infrastructure must be ready whether or not there is device imaging.